OpenStudy (anonymous):

How would you solve 17 --- + 6 ? x+6 This was the original problem: Preform the addition. Simplify if possible. 17 6 ----------- + --------- x^2+12x+36 x+6 I hope I have it right so far :S

OpenStudy (anonymous):

Alg 2?

OpenStudy (anonymous):

It's a developmental class for Rational Expressions, it's from MyMathLab.com.

OpenStudy (anonymous):

Did you factor?

OpenStudy (anonymous):

I did, I got (x+6)(x+6) for the denominator on the right side.

OpenStudy (anonymous):

you need to multiply the top by (x+6) as well; then you can combine the fractions

OpenStudy (anonymous):

\[\frac{ 17 }{ (x+6)^{2} }+\frac{ 6 }{ x+6 }\] I just factored the x^2 +12x+36 and now you need to have a common denominator since one fraction only has one (x+6) you'll need to times another one thus, 17+6(x+6) then distribute 17+6x+36 then just find x

OpenStudy (anonymous):

It's not an equation, it's just an expression; not solving I don't believe.

OpenStudy (anonymous):

Oh my bad, thought it said solve

OpenStudy (anonymous):

Okay so I have what you got factored out

OpenStudy (anonymous):

after that I get confused on what I have to do

OpenStudy (anonymous):

\[\frac{ 6 }{ x+6 }\times \frac{ x+6 }{ x+6 }\]

OpenStudy (anonymous):

6 over x + 6

OpenStudy (anonymous):

Ohh, I think I went wrong with the six, gonna redo it really quick.

OpenStudy (anonymous):

Awesome, I got it. Thanks.

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

Good luck